40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 17 11:09:43 2009, Koji, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC 
    Reply  Tue Nov 17 13:39:37 2009, Koji, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC 
       Reply  Tue Nov 17 15:23:06 2009, Koji, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC 
          Reply  Tue Nov 17 21:10:30 2009, Koji, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC 
             Reply  Wed Nov 18 11:27:33 2009, Koji, josephb, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC 
                Reply  Wed Nov 18 12:33:30 2009, Koji, josephb, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC Untitled.png
       Reply  Wed Nov 18 16:24:25 2009, pete, Configuration, SUS, ETMY suspension conencted to megatron ADC/DAC 
Message ID: 2285     Entry time: Tue Nov 17 21:10:30 2009     In reply to: 2282     Reply to this: 2290
Author: Koji 
Type: Configuration 
Category: SUS 
Subject: ETMY suspension conencted to megatron ADC/DAC 

Koji, Rana

The megatron DAC was temporaly connected to the suspension electronics for the DAC test. We went down to ETMY as we could not excite the mirror.

The DAC is putting correct voltages to the channels. However, the anti imaging filter test output does not show any signal.
This means something wrong is there in the DAC I/F box or the cables to the AI circuit. We will check those things tomorrow.

The ETMY was restored to the usual configuration.

ELOG V3.1.3-